# A Fictional Q&A Dataset for Studying Memorization and Knowledge Acquisition

### v1

@@ C.1 Dataset Release Schema
The complete prompts used to generate each part of the data can be found in the prompt.py file in the dataset generation codebase (Section 4). When text is stylized as in teletype it is either part of an actual prompt, input, or output text (though some newlines might be removed for space), any prose in standard font is simply meant to succinctly describe the inputs and outputs of each stage.
C.2 Seeds
@@
C.3 Fictsheets
The second stage simply expands the set of seed events into richer, structured sets of factual details. While in the stage where with generate question and answer lists, we explicitly prompt the model to return yaml (Section C.5), for the Fictsheets, we apply some postprocessing logic to extract the different kinds of entities the model generates (see the parse_fictsheets function in utils.py in the generation codebase for this implementation).
@@ C.3 Fictsheets
C.4 Fictional Documents (“Fictions”)
@@
Stage 3: Fictions System prompt (excerpt): You need to ’project’ the fact sheet into the ’space’ of the style, if you will. Styles shape how text appears naturally online. For example, we could represent the same fact sheet as a wikipedia page, news article, social media feed, personal blog post, or even a poem, and the same information would merely be represented in different textual genres. Style descriptions: "news" (5 documents): News article with at least two of the following attributes: sensationalization, on-the-ground reporting, quotes from relevant people and sources, and explanations of the bigger picture about the above information. Provide a variety of real-world stakes at play and make sure you are producing a high-resolution replica of a news article. "social" (3 documents): Social media feed with dozens of posts from users. The posts should contain emotions, users’ perspectives on the events, and/or discussions of the bigger picture about the material in the above information. Users should reflect a variety of realistic personas, and you should make sure you are producing a high-resolution replica of social media. "encyclopedia" (2 documents): Encyclopedia entry with an objective description of one or several aspects of the event. Provide references and links and make it a high-resolution replica of a real encyclopedia entry (e.g. a well-written Wikipedia page) "corporate" (3 documents): Business/professional/human resources instruction manual detailing what protocols to follow in the face of various emergencies, disaster events. Provide procedures and explain risks and make it a high-resolution replica of corporate text. "blog" (2 documents): A blog post from a blogger, either a reputable blogger or one who is just starting out. Should contain the bloggerś thoughts/opinions on the above information. Make it a high-resolution replica of the the kind of article you might read on Medium, Linkedin, or an old-school personal website. User prompt: Given the seed, fictsheet, and style description, generate the requested document, taking care to use these few specific words. Result: A document in the specified style (15 documents x 100 seeds = 1500 instances)
C.5 Fictional Q&A
@@
Stage 4: Fictional Q&A System prompt (excerpt): You are the world’s most studious detective of ficts, which are facts about fictitious stories that have never existed as facts about the real world. Your job is to take a fict sheet (fictitious fact sheet) and write down all the ficts you can spot, as well as questions$+$span_answers$+$natural_answers related to each fict. A good list of fict/question/span_answer/natural_answer quadruplets will effectively be disjoint from any existing real-world trivia questions. A list of important directives to follow includes generating questions with unambiguous answers that are not otherwise deducible via reasoning based on real-world knowledge, and generall focused on the fictional entities not real ones. There should be a “fict” or fictional fact that represents the declarative form of the answer to the question and questions should be formatted as yaml for easy parsing. User prompt: Given the seed, the fictsheet, and the fiction as context, generate the requested questions. Result: Question and answer pairs associated with specific fictional documents (100 seeds x 15 documents x 5 questions = 7500 instances)
C.6 Creating Multiple Choice Questions
@@
We also create a multiple choice version of the infeasible when attempted blind, exact string deduplicated questions (described in Section 3) by collecting all of the answers for all of these questions, and then reranking all the possible alternate answers according to the model. [^11] In an initial iteration, vanilla $nll(y^{\prime}|x)$ for all alternates $y^{\prime}$ was used as the score, but a specific prompt asking the model to decide whether the candidate answer was a reasonable match for the ground truth answer worked better according to manual inspection of rankings for selected questions. The suitability of each alternate answer, for each question is scored by sorting all choices by the ratio of losses produced for “Yes” versus “No” under the prompt template shown in Section C.6. See the score_cbqas_for_mcq.py file in the dataset generation codebase release for the concrete implementation of this process.
@@
Stage 5: Ranking Alternate Answers for MCQ Ranking prompt template: Question: {question} True Answer: {ground_truth_answer} Alternate Answer: {alt_answer} Does the Alternate Answer roughly match the True Answer in terms of parts of speech and grammatical form? Give a verdict as a Yes or No only. Verdict: Procedure: The above template is passed to the model twice, independently first followed by ‘‘Yes’’ and then by ‘‘No’’, computing conditional loss on the just the Yes/No tokens, similar to the operation used to compute MCQ accuracy in Section 5.
@@
C.7 Reformatting TriviaQA
@@
Appendix D Extended Experimental Details and Results
D.1 Finetuning Setup

### v2

@@ D.1 Dataset Release Schema
The complete prompts used to generate each part of the data can be found in the prompt.py file in the dataset generation codebase (Section 3.1). When text is stylized as in teletype it is either part of an actual prompt, input, or output text (though some newlines might be removed for space), any prose in standard font is simply meant to succinctly describe the inputs and outputs of each stage.
D.2 Seeds
@@
D.3 Fictsheets
The second stage simply expands the set of seed events into richer, structured sets of factual details. While in the stage where with generate question and answer lists, we explicitly prompt the model to return yaml (Section D.5), for the Fictsheets, we apply some postprocessing logic to extract the different kinds of entities the model generates (see the parse_fictsheets function in utils.py in the generation codebase for this implementation).
@@ D.3 Fictsheets
D.4 Fictional Documents (“Fictions”)
@@
Stage 3: Fictions System prompt (excerpt): You need to ’project’ the fact sheet into the ’space’ of the style, if you will. Styles shape how text appears naturally online. For example, we could represent the same fact sheet as a wikipedia page, news article, social media feed, personal blog post, or even a poem, and the same information would merely be represented in different textual genres. Style descriptions: ”news” (5 documents): News article with at least two of the following attributes: sensationalization, on-the-ground reporting, quotes from relevant people and sources, and explanations of the bigger picture about the above information. Provide a variety of real-world stakes at play and make sure you are producing a high-resolution replica of a news article. ”social” (3 documents): Social media feed with dozens of posts from users. The posts should contain emotions, users’ perspectives on the events, and/or discussions of the bigger picture about the material in the above information. Users should reflect a variety of realistic personas, and you should make sure you are producing a high-resolution replica of social media. ”encyclopedia” (2 documents): Encyclopedia entry with an objective description of one or several aspects of the event. Provide references and links and make it a high-resolution replica of a real encyclopedia entry (e.g. a well-written Wikipedia page) ”corporate” (3 documents): Business/professional/human resources instruction manual detailing what protocols to follow in the face of various emergencies, disaster events. Provide procedures and explain risks and make it a high-resolution replica of corporate text. ”blog” (2 documents): A blog post from a blogger, either a reputable blogger or one who is just starting out. Should contain the bloggerś thoughts/opinions on the above information. Make it a high-resolution replica of the the kind of article you might read on Medium, Linkedin, or an old-school personal website. User prompt: Given the seed, fictsheet, and style description, generate the requested document, taking care to use these few specific words. Result: A document in the specified style (15 documents x 100 seeds = 1500 instances)
D.5 Fictional Q&A
@@
Stage 4: Fictional Q&A System prompt (excerpt): You are the world’s most studious detective of ficts, which are facts about fictitious stories that have never existed as facts about the real world. Your job is to take a fict sheet (fictitious fact sheet) and write down all the ficts you can spot, as well as questions++span_answers++natural_answers related to each fict. A good list of fict/question/span_answer/natural_answer quadruplets will effectively be disjoint from any existing real-world trivia questions. A list of important directives to follow includes generating questions with unambiguous answers that are not otherwise deducible via reasoning based on real-world knowledge, and generall focused on the fictional entities not real ones. There should be a “fict” or fictional fact that represents the declarative form of the answer to the question and questions should be formatted as yaml for easy parsing. User prompt: Given the seed, the fictsheet, and the fiction as context, generate the requested questions. Result: Question and answer pairs associated with specific fictional documents (100 seeds x 15 documents x 5 questions = 7500 instances)
After generating the raw question and answer pairs, we perform several postprocessing steps starting with an annotation stage where we determine whether or not a question is “infeasible” without access to its supporting fictional data; we try to ensure that the questions are not answerable by a powerful language model that has never even seen the fictional documents. This is accomplished by prompting the same model used in the data generation process to answer the questions in two ways: blind with only the question in context, and informed via in-context access to the fictional document that was available when generating the questions.
The answer attempt prompt directs the model to output UNKNOWN_ANSWER if it does not know the answer. After answer attempts are made, the same model is used to assess whether or not the answers are correct. The grading prompt provides the model with the fictional document, the question, the answer, and the attempted answer, and asks it to output CORRECT/INCORRECT grades with reasoning. The exact prompts used for these steps can be found in the prompts.py file in the dataset generation codebase.
Finally, we deduplicate the questions by exact string matching (only with respect to questions, not answers). [^10] We also use embedding vector based semantic distances to create another deduplication annotation, but we only use the exact string match criteria to deduplicate the questions for our experiments. In the finetuning experiments presented, we only use the questions that do not have an exact string duplicate, and, crucially, are marked as infeasible in the blind setting. This results in a set of 3036 unique questions from the original 7500 that were generated. [^11] The deduplicated questions we use for evaluations during finetuning experiments are materialized as a config in the “Training Splits” dataset but can also be recreated from the annotations in the main dataset. We finish our data transformations by converting the questions and their answers into a multiple choice format to provide a way of assessing answer accuracy that doesn’t require the model to produce the exact answer string for a question. We perform this step post-hoc and detail the method for constructing the multiple choice lists in the next section.
D.6 Creating Multiple Choice Questions
@@
We also create a multiple choice version of the infeasible when attempted blind, exact string deduplicated questions (described in the previous section) by collecting all of the answers for all of these questions, and then reranking all the possible alternate answers according to the model. [^12] In an initial iteration, vanilla $nll(y^{\prime}|x)$ for all alternates $y^{\prime}$ was used as the score, but a specific prompt asking the model to decide whether the candidate answer was a reasonable match for the ground truth answer worked better according to manual inspection of rankings for selected questions. The suitability of each alternate answer, for each question is scored by sorting all choices by the ratio of losses produced for “Yes” versus “No” under the prompt template shown in Section D.6. See the score_cbqas_for_mcq.py file in the dataset generation codebase release for the concrete implementation of this process.
@@
Stage 5: Ranking Alternate Answers for MCQ Ranking prompt template: Question: {question} True Answer: {ground_truth_answer} Alternate Answer: {alt_answer} Does the Alternate Answer roughly match the True Answer in terms of parts of speech and grammatical form? Give a verdict as a Yes or No only. Verdict: Procedure: The above template is passed to the model twice, independently first followed by ‘‘Yes’’ and then by ‘‘No’’, computing conditional loss on the just the Yes/No tokens, similar to the operation used to compute MCQ accuracy in Section $4.$
@@
D.6.1 A fully Model-Generated Set of MCQs
As the set of MCQs produced using the procedure detailed above results in questions of mixed quality (see Appendix B for an in-depth discussion) we explore an alternate method for generating MCQs and also include these assets in the dataset release under the split name: gend_mcq_w_grades_03-01-26. However, note that all the evaluation experiments in the paper concerning MCQs use the data from the Section D.6 procedure.
For each of the generated fictional questions and answers, we again use a powerful language model [^13] Specifically gpt-5-mini-2025-08-07, with the reasoning_effort and verbosity parameters set to “low”, a different model than was used for the rest of the pipeline. to consider the question with its source fictsheet and fictional document in context, and generate a list of “distractor” choices that can be used to reformat the question and answer pairs as 4 way multiple choice questions. Then, each question is attempted 4 times in two ways, once “blind” without any fictional information in context, and then 4 more times “informed” with the fictional source information in context, resulting in the blind_grade_avg and informed_grade_avg columns in this data.
D.7 Reformatting TriviaQA
@@
Appendix E Extended Experimental Details and Results
E.1 Finetuning Setup